• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Higinbotham

    Log entry time 16:35:19 on August24,2009

    Entry number 285712

    keyword=Shift Summary

    0900 Hall in Controlled Access for Target Work
    1000 Spectrometers Set to Zero Momentum for Target Cool Down
    1230 Target Window Back In Place
    1315 No Cool Down Today, Setting Spectrometers Back to working momentum
    1600 Jack and Dave back into the Hall (Jack is looking at right dipole)